OpenStudy (anonymous):

The mass of Neptune is about 1.03 x 10^23 metric tons. The mass of the sun is about 1.998 x 10^27 metric tons. About how many times the mass of Neptune is the mass of the sun. Give the answer in scientific notation.

OpenStudy (saifoo.khan):

The mass of neptune is about 1.03x10^23 metric tons. The mass of the sun is about 1.998x10^27 metric tons. ABout how many times the mass of Neptune is the mass of the sun? (give your answer in scientific notation) Mass of sun = K*(mass of neptune) K=mass of sun/mass of neptune = 1.998x10^27/1.03x10^23 = 1.939*10^4

OpenStudy (anonymous):

Hi Saifoo!! It says that's not correct...also says "divide carefully, and give your answer in the form M x 10^n, where M is greater than or equal to 1 and less than 10 (whatever that means).

OpenStudy (saifoo.khan):

HI bunny, Oh. ok, let me check..

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

oh, it also says to round to the nearest thousandth

OpenStudy (saifoo.khan):

I remember this one. I don't remember this answer. Can we skip this? I can help you asking a friend of mine.

OpenStudy (anonymous):

okay, thank you. I'll wait and go to another.

OpenStudy (saifoo.khan):

hmm, :D